### v1

# Free Record-Level Privacy Risk Evaluation Through Artifact-Based Methods

Abstract
Membership inference attacks (MIAs) are widely used to empirically assess the privacy risks of samples used to train a target machine learning model. State-of-the-art methods however require training hundreds of shadow models, with the same size and architecture of the target model, solely to evaluate the privacy risk. While one might be able to afford this for small models, the cost often becomes prohibitive for medium and large models.
We here instead propose a novel approach to identify the at-risk samples using only artifacts available during training, with little to no additional computational overhead. Our method analyzes individual per-sample loss traces and uses them to identify the vulnerable data samples. We demonstrate the effectiveness of our artifact-based approach through experiments on the CIFAR10 dataset, showing high precision in identifying vulnerable samples as determined by a SOTA shadow model-based MIA (LiRA [7]). Impressively, our method reaches the same precision as another SOTA MIA when measured against LiRA, despite it being orders of magnitude cheaper. We then show LT-IQR to outperform alternative loss aggregation methods, perform ablation studies on hyperparameters, and validate the robustness of our method to the target metric. Finally, we study the evolution of the vulnerability score distribution throughout training as a metric for model-level risk assessment.
**footnotetext: Equal contribution
1 Introduction
Figure 1: Precision@k=1% (250 samples) when identifying vulnerable samples determined by LiRA attack at a variable FPR threshold (log scale).
Machine Learning (ML) models are increasingly trained or fine-tuned by companies, governments, and academic researchers often using private, sensitive, or confidential data. This includes medical [17], legal [11], and financial data [3].
Membership inference attacks (MIAs) are widely used to empirically evaluate whether an ML model has memorized individual records and the overall privacy risk releasing the model poses. ML models have indeed been shown to memorize their training data [35, 8, 7], allowing an adversary to detect the presence of a certain data sample in the training dataset [41, 34, 7], or even reconstruct the entire training samples with little prior knowledge [6, 14].
Model developers often rely on running empirical privacy attacks against their own models to assess potential risks associated with the public release of these models or their predictions [31]. This empirical approach to privacy analysis is often used alongside the privacy-preserving training methods with formal guarantees, such as DP-SGD [1], or as a standalone mitigation where the severe impact of differentially private training on model utility makes it impractical [31].
State-of-the-art privacy attacks are, however, computationally expensive to run. They require training multiple, often hundreds, of shadow models [7, 40, 45], with each shadow model requiring the same amount of computational resources as one original training run. Depending on the scale of the target model and the training dataset, this renders practical uses of such attacks challenging–e.g. as part of the selection process of architectures, training parameters, etc–if not completely infeasible - for example, in the case of Large Language Models.
Contribution. We here propose a new class of methods to estimate a model’s vulnerability to privacy attacks exclusively utilizing artifacts available for free at training time. Artifact-based methods allow model developers to leverage their white-box access to the model throughout the training to identify data samples vulnerable to privacy attacks with little to no computational overhead.
To first underscore the difficulty of the task, we study the level of agreement between SOTA MIA methods and show them to only partially agree on the most vulnerable points. Despite assuming the same attacker knowledge and being similar shadow model-based methods, LiRA and Attack R only partially agree on the most vulnerable points.
We then consider a binary classification setup, defining vulnerable samples as the ones successfully identified as members by the Likelihood Ratio Attack (LiRA) attack [7] at a fixed low False Positive Rate (FPR), a highly imbalanced task.
We then propose Loss Trace IQR (LT-IQR), as a simple yet robust aggregation score for sample-level loss. In a series of experiments on the CIFAR10 [19] dataset, we confirm our proposed method to reach a level of agreement on par with computationally expensive SOTA shadow model-based MIAs despite being essentially free. LT-IQR (75%-25%), for instance, achieves 61% Precision@k=1% in predicting vulnerable samples defined by LiRA at a low FPR threshold of $10^{-3}$, a result in line with Attack R [40]. This demonstrates that our method is as good at identifying vulnerable samples as a state-of-the-art shadow-model-based MIA, while being essentially free for the model developer.
Additionally, we compare LT-IQR to multiple potential aggregation methods for the loss trace and show it to outperform these. We perform ablation studies on the choice of quantile parameters in LT-IQR, and validate the robustness of our method to the target metric. Finally, we study the evolution of the vulnerability score distribution across the training process as a metric for model-level risk assessment.
Our cheap yet effective approach enables model developers to assess the privacy risks associated with a given model release, and deploy appropriate mitigations if needed, including unlearning, data removal, or retraining with differential privacy. Given its low computational cost, our method could be especially useful to compare the privacy leakage of different hyperparameter setups, or as an early indication to run a full privacy attack on a subset of candidate models.
2 Related Work
2.1 Membership Inference Attacks
Membership Inference Attacks (MIAs) aim to determine whether a certain data sample was used to train a given model. They typically rely on a strong yet realistic attacker [31, 28] and are used as an upper bound for other attacks including extraction  [10, 30, 25] and attribute inference attacks.
First introduced in the context of ML training by Shokri et al. [34], current MIAs against ML models typically assume a black-box setting and rely on the loss of the trained model as the main signal for memorization [34, 41, 7, 40, 45]. Sablayrolles et al. [33], for instance, argue that an optimal attack only needs to depend on the loss and that white-box access would give no additional advantage to an attacker.
Widely adopted MIAs today, including LiRA [7] and Attack R [40], however involve training a large number of shadow models [34, 7, 24, 40]. This makes their practical application very expensive, and in many cases infeasible.
While attacks have been developed with lower computational costs [41, 45], they either suffer from low utility compared to the shadow-model-based attacks [41], or still require training some shadow models [45], an expensive task, with the attack performance being positively correlated with the number of shadow models.
2.2 Individual vulnerability
Thanks to the extensive literature on the topic, vulnerability to privacy attacks is now understood to be highly non-uniform across data distributions and is measured via TPR@low FPR. Empirical privacy risk is indeed typically concentrated around a small number of data outliers [13, 5, 9]. Additionally, recent works [2, 15] argued that common MIA evaluations fail to capture the risk for the most vulnerable samples. Throughout this work we focus on record-specific risk, with the main goal to identify the samples most vulnerable to privacy attacks.
While most MIAs are postulated as a privacy game with a fixed model [34, 41, 7], Ye et al. [40] have adapted the definition to use a fixed record instead, focusing on individual vulnerability. Subsequent work has developed record-specific attacks [4] and evaluated existing techniques under the new definition [15, 20]. In particular, Leemann et al. [20] investigate the connection between MIA vulnerability and certain per-instance properties: final loss, parameter gradients, local loss Hessian, and others, identifying the final loss to be the best predictor of the MIA performance. Additionally, Tobaben et al. [38] study the impact of the average gradient norm throughout training on the MIA performance and find a strong correlation across multiple datasets. Unlike our work, however, both Leemann et al. [20] and Tobaben et al. [38] study the correlation between the data properties and attack success while we focus on using training artifacts to accurately identify the few most vulnerable samples.
In the context of differentially private training [12, 1], individual data-dependent privacy accounting has been proposed, recognizing non-uniform privacy leakage across training data samples [42, 18, 37]. Unlike classic data-independent privacy accounting mechanisms [1, 26, 27], individual privacy accounting considers per-sample gradient norms and yields stronger privacy guarantees for data points falling below the clipping threshold. As such, Thudi et al. [37] show that on common benchmark datasets, many individual data samples enjoy significantly better privacy protection than what is suggested by the conservative data-independent privacy analysis traditionally used for DP-SGD.
2.3 Tracing memorization
Unlike privacy analyses described above, we consider an artifact-based approach where an attacker, here the model developer, has access to all intermediate model states during training. This setup, for instance, allows us to use the loss throughout training as an additional source of information on top of the commonly used final loss.
Some prior works have considered tracing privacy metrics throughout training. Pruthi et al. [32] tracks the loss before and after each training step to estimate the influence of a training sample on model’s predictions. Lesci et al. [21] characterize a model’s counterfactual memorization [46] by observing losses on a subset of data throughout training. Liu et al. [23] and Li et al. [22] propose using loss trajectory in a traditional MIA setup by tracking the loss during the target model distillation process.
For differentially private training, privacy auditing often adopts a similar threat model to us. Privacy auditing has emerged as a way to either assess the tightness of the theoretical privacy analysis [16, 30, 44] or verify the correctness of DP algorithm implementations [39, 29]. Privacy auditing empirically estimates the lower bound for the differential privacy guarantees of the model training process by running privacy attacks and measuring their success. Assumed threat models can range from a weak adversary with only black-box access to the model [16, 30, 36], to a white-box adversary capable of observing intermediate model updates [30, 29, 36], to a very strong adversary with full control over the training dataset apart from the target record [30]. Prior works on privacy auditing however typically rely on data poisoning attacks, instantiating the MIAs against data samples (canaries) purposely injected into the training dataset.
Proposed attack methods vary depending on the threat model. For the black-box attacks, a simple loss threshold [41] is typically used. For white-box adversaries with access to the intermediate model states, a threat model similar to the one we assume in this work, prior works suggested using aggregate metrics over the entire training process. Proposed methods include taking the maximum or average of the loss [30], or computing the dot product between the canary gradient and the overall batch update [29, 36].
3 Preliminaries
3.1 Problem setup
3.1.1 Model training
Let $f_{\theta}:\mathcal{X}\rightarrow[0,1]^{N}$ be a machine learning classification model with an input domain $\mathcal{X}$ and an output domain $\mathcal{Y}$ with $N=|\mathcal{Y}|$ classes, parametrized by a parameter vector $\theta$. $f_{\theta}(x)$ outputs a probability distribution over $N$ classes. For brevity, we denote the probability of a given class $y$ as $f_{\theta}(x)_{y}$. We denote train and test datasets as $D_{\text{train}}=\{(x_{i},y_{i})\mid x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y}\}$ and $D_{\text{test}}=\{(x^{*}_{i},y^{*}_{i})\mid x^{*}_{i}\in\mathcal{X},y^{*}_{i}\in\mathcal{Y}\}$ respectively. Let $\mathcal{T}$ be a training process such that the final model weights $\theta$ are obtained as $\theta\leftarrow\mathcal{T}(D_{\text{train}})$. We consider the training process to be performed over $S$ epochs, and additionally denote intermediate model state after $s=1\dots S$ epochs as $\theta_{s}\leftarrow\mathcal{T}_{s}(D_{\text{train}})$. Note that $\theta_{0}$ is the initial model state and $\theta_{S}=\theta$ is the parameter vector of a fully trained model.
We consider training with cross-entropy loss defined as:
$\ell(f_{\theta}(x),y)=-\log{f_{\theta}(x)_{y}}$ (1)
We denote the loss on a sample $(x,y)$ after $s$ training epochs as $\ell_{s}(x,y)=\ell(f_{\theta_{s}}(x),y)$. We also define $\mathcal{L}_{\text{train}}^{s}$ and $\mathcal{L}_{\text{test}}^{s}$ as average losses at epoch $s$ on train and test sets respectively.
$\mathcal{L}_{\text{train}}^{s}=\frac{1}{|D_{\text{train}}|}\sum_{(x,y)\in D_{\text{train}}}\ell_{s}(x,y)$ (2)
$\mathcal{L}_{\text{test}}^{s}=\frac{1}{|D_{\text{test}}|}\sum_{(x^{*},y^{*})\in D_{\text{test}}}\ell_{s}(x^{*},y^{*})$ (3)
Training is performed in mini-batches $\mathcal{B}_{i}$ with each optimization step defined as
$\theta^{t+1}\leftarrow\theta^{t}-\eta\sum_{(x,y)\in\mathcal{B}_{i}}\nabla_{\theta}\ell(f_{\theta}(x),y)$ (4)
where $\tilde{\theta}$ is the updated parameter vector and $\eta$ is the learning rate.
3.1.2 Membership Inference Attacks
A membership inference attack (MIA) aims to determine whether a given data sample was used to train a model. Formally, we define a membership inference attack as a function $A:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow[0,1]$, where $A(x,y,\theta)=1$ indicates that the attacker believes the sample $(x,y)$ was used to train the model with parameters $\theta$, and $A(x,y,\theta)=0$ indicates when they believe it was not. In particular, we consider SOTA MIAs that output a confidence score, where higher values indicate stronger belief that the sample was used during training. The binary classification decision is then made by comparing this score against a threshold $\tau$:
$A_{\tau}(x,y,\theta)=\begin{cases}1&\text{if }A(x,y,\theta)\geq\tau\\ 0&\text{otherwise}\end{cases}$ (5)
For evaluation purposes, we consider a balanced dataset $D=D^{+}\cup D^{-}$ where $D^{+}\subseteq D_{\text{train}}$ represents member samples and $D^{-}\subseteq D_{\text{test}}$ represents non-member samples, with $|D^{+}|=|D^{-}|$. The attack’s performance is measured using standard binary classification metrics:
  • True Positive Rate (TPR): The fraction of member samples correctly identified as members
    $\text{TPR}(\tau)=\frac{|\{(x,y)\in D^{+}:A(x,y,\theta)\geq\tau\}|}{|D^{+}|}$ (6)
  • False Positive Rate (FPR): The fraction of non-member samples incorrectly identified as members
    $\text{FPR}(\tau)=\frac{|\{(x,y)\in D^{-}:A(x,y,\theta)\geq\tau\}|}{|D^{-}|}$ (7)
Throughout this work, we focus on identifying the most vulnerable samples by fixing a low FPR threshold $\alpha$ (typically $\alpha=10^{-3}$) and considering samples that are identified as members at this strict threshold by SOTA shadow model-based MIAs. We use the terms “point” and “sample” interchangeably. For a given $\alpha$, we select the threshold $\tau_{\alpha}$ such that $\text{FPR}(\tau_{\alpha})=\alpha$.
Definition 3.1 (Vulnerable points).
We define the set of vulnerable points $V$ at a given FPR threshold $\alpha$ as:
$V_{\alpha}=\{(x,y)\in D_{\text{train}}:A(x,y,\theta)\geq\tau_{\alpha}\}$ (8)
This allows us to evaluate the precision of our vulnerability detection method by comparing against the set of vulnerable points identified by SOTA shadow model-based (SM-based) MIA techniques. The focus on vulnerable samples identified at a strict FPR threshold $\alpha$ captures the samples that are most clearly distinguishable as training set members. These points represent the highest privacy risk, as they can be confidently identified by SOTA SM-based MIAs as members even under conservative attack settings that minimize false positives. By fixing a low FPR threshold (e.g., $\alpha=10^{-3}$), we ensure that any point classified as vulnerable represents a genuine privacy concern rather than a statistical artifact, since the attacker must be highly selective to maintain such a low false positive rate.
3.1.3 Likelihood Ratio Attack (LiRA)
LiRA [7] frames membership inference as a likelihood ratio test between two hypotheses. For a target point $(x,y)$, it compares the likelihood of observing the model’s loss under the hypothesis that $(x,y)$ was a training set member versus a non-member. Specifically, it estimates:
$A_{\text{LiRA}}(x,y,\theta)=\log\frac{p(\ell(f_{\theta}(x),y)|(x,y)\in D_{\text{train}})}{p(\ell(f_{\theta}(x),y)|(x,y)\not\in D_{\text{train}})}$ (9)
To estimate these likelihood distributions, LiRA trains $N$ pairs of shadow models $\{(M_{i}^{+},M_{i}^{-})\}_{i=1}^{N}$, where $M_{i}^{+}$ includes $(x,y)$ in its training set and $M_{i}^{-}$ excludes it. The shadow models provide samples from the loss distributions conditioned on membership and non-membership, allowing estimation of the likelihood ratio.
3.1.4 Attack R
Attack R [40] frames membership inference as a sample-dependent test by comparing the target model’s loss against reference models trained without the target point. For a target point $(x,y)$, it computes:
$A_{\text{Attack R}}(x,y,\theta)=\text{Pr}[\ell(f_{\theta}(x),y)\leq c_{\alpha}(x,y)]$ (10)
where $c_{\alpha}(x,y)$ is a threshold that depends on the specific target point $(x,y)$. To estimate this threshold, Attack R trains $N$ reference models $\{M_{i}\}_{i=1}^{N}$ on randomly sampled datasets that exclude $(x,y)$. The threshold $c_{\alpha}(x,y)$ is set as the $\alpha$-percentile of the empirical loss distribution $\{\ell(f_{M_{i}}(x),y)\}_{i=1}^{N}$ to achieve a target false positive rate $\alpha$.
Contrary to LiRA, Attack R only trains reference models without the target point (rather than pairs with/without), making it more computationally efficient.
4 Method
(a) Loss traces of 3 samples from the “frog” class in CIFAR-10 representing different example hardness.
(b) Average loss traces of 2 groups of 100 points representing different example hardness and the average proportion of vulnerable points per group, compared to the average test loss.
Figure 2: Plots illustrating the intuition behind loss traces from CIFAR-10 dataset
4.1 Intuition
The main challenge of running loss-based MIAs stems from the fact that the final loss for a given sample is influenced by two independent factors: example hardness and memorization [7]. For instance, extremely low loss values can either represent a very easy-to-classify non-member example or a well-memorized member. Similarly, a relatively high loss value can represent either a hard-to-classify, partially memorized member, or an average non-member. Identifying whether a given sample is easy- or hard-to-classify is therefore a key component of a strong loss-based MIA.
MIAs relying on shadow models [7, 40] account for example hardness by training a number of models on datasets without the target sample. The resulting loss distributions of the target sample are then measured, and used to inform the attack.
We here posit that example hardness might also be inferred, for free, from training artifacts such as the loss trace over training. The intuition is that easy-to-fit samples would exhibit an early drop in the training loss and quickly reach a small loss value. Meanwhile, hard-to-classify yet memorized samples would take more epochs to converge to a final, and equally small, loss value. A similar logic applies to samples with moderately high losses at the end of training. If they are very hard-to-classify samples exhibiting even higher losses on non-member shadow models, they could still be correctly identified by an MIA. Here, a steep drop in loss during training (even if not reaching zero) would signal that a sample is partially memorized.
Fig. 2(a) shows loss traces on three samples from the CIFAR-10 “frog” class, collected during training (all samples were part of the training dataset). These traces represent distinct types of loss behavior. The purple sample is a typical sample for the frog class: it is easily learned by the model and the train loss quickly converges to small loss values. The olive sample, on the other hand, is an outlier frog sample. It features unique characteristics compared to other samples from the same class, notably the color. Here we can see that the loss is quite noisy and remains fairly high until approximately epoch 70 when it gets memorized. Finally, the cyan sample is an even harder-to-predict sample from the frog class. The early loss (around epoch 10) for this sample is significantly higher than for the purple example. The loss then decreases, but does not reach the low values exhibited by the purple and olive samples by the end of training.
Following this intuition, we further verify our hypothesis by grouping samples with similar loss traces together, and measuring the proportion of samples considered vulnerable by LiRA [7] at $\alpha=10^{-3}$ (see Def. 3.1) in these groups. In particular, on Figure 2(b) we consider average loss traces for two groups with 100 samples each, alongside the average train loss. Both groups are selected to have relatively high loss at an early sample in the training (epoch 11, vertical dashed line on the graph) making them more likely to be outliers and at-risk samples. One group then converges to near-zero training loss, representing easy-to-fit outliers. The final loss for the final group remains relatively high, thus representing hard-to-fit outliers. Borrowing the nomenclature from Carlini et al. [7], we expect a) outliers to be more vulnerable to MIAs and b) easy-to-fit samples to be more vulnerable than hard-to-fit samples. Figure 2(b) indeed confirms that vulnerable samples make up 29% of the easy-to-fit outlier group compared to 12% for hard-to-fit outliers. When comparing to the average baseline of 5%, we confirm that samples with high loss at the early stage of training, outliers, are more vulnerable to MIAs and that, among outliers, easy-to-fit samples (defined by the low final loss) are more vulnerable than hard-to-fit ones. Our method aims to utilize this information to identify at-risk samples.
4.2 Computing per-sample losses
In most cases, computing the per-sample loss during training incurs little overhead in addition to normal training. For instance, we consider the typical PyTorch training loop below, where outputs and targets represent batched data.
criterion = nn.CrossEntropyLoss()
loss = criterion(outputs, targets)
# here the loss is averaged out and
# does not have a batch dimension
loss.backward()
Here the loss reduction is set to mean by default and the loss is averaged over a batch before being returned from the criterion call. Instead, we can easily perform the aggregation on the client side, storing the per-sample losses before averaging.
criterion = nn.CrossEntropyLoss(
reduction="none"
loss = criterion(outputs, targets)
# here the loss is per-sample loss
# and does have a batch dimension
saved_losses.append(
loss.detach()
# take the mean before backward pass
loss.mean().backward()
We note, however, that this approach is not applicable if augmentations are used. In this case tracking per-epoch losses incurs the cost of one forward pass on the full dataset on every epoch.
Memorization is indeed a property of a specific sample: the model’s behaviour on memorized samples and their augmented versions might typically be very different. As such, when augmentations are applied during training, losses computed during training would be losses on augmented samples, not the original ones. In this scenario we suggest recording the losses of the non-augmented training samples in a separate forward pass after each epoch. While slightly more costly, this approach remains computationally much cheaper than training even one shadow model, as a forward pass is typically several times cheaper than the backward pass.
4.3 Loss trace aggregation
For each training sample $(x,y)\in D_{\text{train}}$ we consider a sequence of per-epoch losses (See Sec. 3.1.1):
$\text{trace}(x,y)=[\ell_{0}(x,y),\ell_{1}(x,y),\dots,\ell_{S}(x,y)]$
where $S$ is the total number of training epochs.
To make use of the available training artifacts and keep the additional computations to a minimum, we approximate $\ell_{i}(x,y)$ (defined in Sec. 3.1.1 as the loss after the $i$-th epoch) as the training loss computed during the $i$-th epoch. We then consider multiple aggregating functions to compute the vulnerability score $\phi$.
Note that, where relevant, early epoch $s^{*}$ is defined as an epoch at which the average test loss $\mathcal{L}_{\text{test}}^{s^{*}}$ has decreased significantly compared to the initial $\mathcal{L}_{\text{test}}^{0}$, but the train/test gap ($\mathcal{L}_{\text{test}}^{s^{*}}-\mathcal{L}_{\text{train}}^{s^{*}}$) remains low. Intuitively, we consider epoch $s^{*}$ as a point where the model has gained some generalization ability, but has not significantly memorized any samples.
Below we discuss various aggregation methods on the loss trace.
Final loss
This represents the final model loss at epoch $S$.
$\phi(x,y)=\ell_{S}(x,y)$ (11)
Mean loss
$\phi(x,y)=\frac{1}{S}\sum_{s=1}^{S}\ell_{s}(x,y)$ (12)
Loss delta
This represents how much the model has learned about this particular point.
$\phi(x,y)=\ell_{s^{*}}(x,y)-\ell_{S}(x,y)$ (13)
Smooth loss delta
Intuitively, this represents the same property as loss delta, but the values are averaged over multiple epochs (for a certain value of $\delta$) for robustness.
$\phi(x,y)=\frac{1}{2\delta}\sum_{s=s^{*}-\delta}^{s^{*}+\delta}\ell_{s}(x,y)-\frac{1}{2\delta}\sum_{s=S-2\delta}^{S}\ell_{s}(x,y)$ (14)
Normalized loss delta
It measures the fraction of a sample’s early-stage loss that was eliminated by the end of training, capturing the relative (rather than absolute) improvement in the model’s performance on that sample.
$\phi(x,y)=\frac{\ell_{s^{*}}(x,y)-\ell_{S}(x,y)}{\ell_{s^{*}}(x,y)}$ (15)
Loss Trace IQR (LT-IQR)
This robustly measures how much a sample’s loss fluctuates during training by computing the spread between two chosen quantiles of its loss trace, capturing the volatility in how well the model learns the specific example over time. Below $Q_{q_{1}}$ and $Q_{q_{2}}$ are the $q_{1}$-th and $q_{2}$-th quantiles of the loss values across all epochs, with $0\leq q_{1}<q_{2}\leq 1$
$\phi(x,y)=Q_{q_{2}}(\{\ell_{s}(x,y)\}_{s=1}^{S})-Q_{q_{1}}(\{\ell_{s}(x,y)\}_{s=1}^{S})$ (16)
5 Evaluation
5.1 Experimental setup
The code to reproduce the experiments is available at https://github.com/computationalprivacy/loss_traces.
5.1.1 Target model
Adopting the evaluation setup from the literature [7], we train Wide ResNet28-2 [43] models on the CIFAR-10 dataset [19] for 100 epochs, reaching a test accuracy of $\sim$86%. We train using a stochastic gradient descent optimizer, with a momentum of 0.9 and weight decay set to 0.0001. We use a cosine annealing learning rate scheduler with initial value of 0.1. We use a batch size of 256. During training, we augment the images via random horizontal flips, and random crops with four pixels of padding, following the augmentation setup used for training by Carlini et al. [7].
We randomly split the training part of CIFAR-10 into member and non-member subsets, each containing 25000 samples. We then use the member subset for training.
We track the per-sample losses every epoch. As we use augmentations, we collect training losses on the non-augmented samples during the evaluation loop after each epoch.
Where appropriate, we set the early stage epoch $s^{*}$ (see Sec. 4.3) at 11. At this stage the average test loss $\mathcal{L}_{\text{test}}^{s^{*}}$ has decreased by $\sim 75$%, while the train/test gap $\mathcal{L}_{\text{test}}^{s^{*}}-\mathcal{L}_{\text{train}}^{s^{*}}$ remains low at $\sim 2$%.
5.1.2 Membership Inference Attacks
We evaluate three distinct membership inference attacks: LiRA, Attack R, and LOSS. We follow the setup proposed in Carlini et al. [7], instantiating online LiRA with 256 shadow models (128 member and 128 non-member models respectively). We then also instantiate Attack R with 128 non-member shadow models. All shadow models are trained using the same training procedure described for the target model in Sec. 5.1.1, with each shadow model trained on a different member/non-member split. For LiRA, we use 2 queries per sample through horizontal flipping, matching the “mirror augmentation” setup from Carlini et al. [7].
Figure 3: ROC curves for LiRA, Attack R, and LOSS membership inference attacks on CIFAR-10 against Wide ResNet28-2 (log scale).
MIA AUC TPR@FPR=$10^{-3}$
LiRA [7] $0.71$ $0.049$
Attack R [40] $0.64$ $0.013$
LOSS [41] $0.57$ $0.001$
Table 1: AUC and TPR@FPR=$10^{-3}$ scores for LiRA, Attack R and LOSS MIAs with our training setup.
Table 1 shows the performance of MIAs we consider in our training setup. We evaluate them on a random balanced split of members and non-members from the CIFAR-10 train split. As expected, SOTA shadow model-based attacks (LiRA, Attack R) significantly outperform LOSS, with LiRA showing the best performance. Note however that though the LOSS attack shows non-trivial AUC=$0.57$, it fails to confidently identify members - as measured by TPR@FPR=$10^{-3}$ being at random chance. Additionally, Fig. 3 shows full AUC curves for the three attacks.
5.1.3 Evaluation metric
We evaluate our vulnerability detection methods using Precision@k and Recall@k, which measures what fraction of the k points with highest vulnerability scores are actually vulnerable according to a reference MIA (LiRA). For a vulnerability scoring function $\phi$ and a set of vulnerable points $V_{\alpha}$ (for a given FPR threshold $\alpha$ as defined in Definition 3.1), we define Precision@k as:
[…]
where $\text{top}_{k}(\phi)$ represents the k points with the highest vulnerability scores according to $\phi$. In other words, if we sort all training points by their vulnerability score $\phi(x,y)$ in descending order and take the first k points, Precision@k tells us what fraction of these points are actually vulnerable according to our reference MIA.
Similarly, we define Recall@k as:
[…]
Recall here measures the fraction of all vulnerable points we successfully identified within our top k predictions.
The results we report are averaged across 10 models trained with varying randomization seed. Throughout the paper, we use LiRA as the reference MIA.
5.2 Cross-MIA correlation
Figure 4: Confusion matrix comparing the predictions of vulnerable and non-vulnerable points identified by LiRA and Attack R where we consider the top 1000 points to be vulnerable.
Before evaluating our method’s effectiveness at identifying vulnerable points, it’s important to understand the inherent challenges of this task.
First, vulnerable points are rare: at a strict false positive rate of $10^{-3}$, LiRA, for instance, identifies only 1,153 samples to be at-risk, slightly below 5% of our training set of 25,000 samples. This means that finding vulnerable points is a significantly more challenging task than for instance the standard member vs. non-member classification. Indeed, even small absolute errors can lead to large relative performance drops when dealing with such a small target set and a random guess baseline for this task would be below 5%.
Furthermore, we study the level of agreement between SOTA methods on which points are the most at-risk and show, somewhat surprisingly, that there is only a partial agreement on which points are most vulnerable. Figure 4 shows the overlap between the top 1,000 most vulnerable points identified by each attack. Despite both LiRA and Attack R assuming the same attacker, being both shadow model-based and part of the same class of attacks, and achieving strong overall membership inference performance (AUC of 0.71 and 0.64 respectively, see Table 1), they only share 551 points among their top 1,000 predictions. We believe this to be both an important yet understudied aspect of MIAs, with potentially important practical consequences.
k = 1% k = 3% k = 5%
Method Precision@k Recall@k Precision@k Recall@k Precision@k Recall@k
Attack R [40] $0.62$ $0.13$ $0.60$ $0.37$ $0.57$ $0.59$
LOSS [41] $3\times 10^{-4}$ $7.2\times 10^{-5}$ $6.7\times 10^{-4}$ $3.7\times 10^{-4}$ $5.6\times 10^{-4}$ $5.4\times 10^{-4}$
Final loss $0.08$ $0.02$ $0.11$ $0.07$ $0.12$ $0.12$
Mean loss $0.36$ $0.07$ $0.38$ $0.24$ $0.36$ $0.37$
Loss delta $0.09$ $0.02$ $0.10$ $0.06$ $0.07$ $0.08$
Normalized loss delta $0.14$ $0.03$ $0.12$ $0.07$ $0.12$ $0.12$
LT-IQR (75%-25%) $0.61$ $0.13$ $0.49$ $0.30$ $0.42$ $0.43$
Table 2: Precision@k and Recall@k for various loss trace aggregation methods, reported at fixed FPR=$10^{-3}$ and varying values of $k$. We also report the performance of existing MIAs on the same task for the reference
Figure 5: Recall@k=1% (250 points) for identifying vulnerable points determined by LiRA attack at a variable FPR threshold (log scale).
5.3 LT-IQR empirical results
We report the precision of different vulnerability detection methods (Final loss, Attack R, us) and random guess at identifying points that are the most vulnerable, defined as LiRA at low FPR points.
Figure 1 shows that our method succeeds at identifying the most vulnerable points (a low FPR regime) according to LiRA, and does so as well as Attack-R, while being orders of magnitude computationally cheaper. Focusing on the FPR=$10^{-3}$ threshold, which corresponds to a highly selective attack setting, our best performing aggregation metric (LT-IQR (75%-25%)) achieves approximately 60% precision at $k=1\%$ of training data. This means that among the 250 training points our method identifies as most vulnerable, 61% are actually vulnerable according to LiRA at this strict false positive threshold. This, impressively, matches the performance of Attack R despite requiring no shadow model training compared to Attack R’s 128 shadow models. It also significantly outperforms simpler baselines like final loss (8% precision) and random guess.
Looking at the broader trend across FPR thresholds, we observe that LT-IQR’s performance advantage does persist. At stricter thresholds, LT-IQR slightly outperforms Attack R while strongly outperforming both final loss and random baselines. As we relax the FPR threshold (moving right on the graph), Attack R starts to show slightly higher precision compared to our method for medium-vulnerability points.
We then consider the recall of the predictions. Figure 5 evaluates the recall of different vulnerability detection methods across varying FPR thresholds, examining what fraction of truly vulnerable points each method successfully identifies within its top k=1% predictions. The results reveal several key patterns. First, at very strict thresholds ($\text{FPR}<10^{-3}$), LT-IQR identifies around 50% of vulnerable points in its highest-ranked predictions - meaning we capture half of these extremely vulnerable samples by predicting just 250 samples (1% of the training data). As we relax the FPR threshold, recall will then naturally decrease: while more points become classified as vulnerable at higher FPR thresholds, our method still examines only the same top $k=250$ predictions. The fact that LT-IQR maintains higher recall than Attack R until FPR=$10^{-3}$ (and closely matches it for higher FPR values), demonstrates its robustness as a sample vulnerability detection method. At more permissive thresholds ($\text{FPR}>0.01$), all methods, including random guessing, begin to achieve similar recall, reflecting the fact that the task becomes less discriminative when a large fraction of points are considered vulnerable.
Taken together, these results demonstrate that our simple, artifact-based LT-IQR method succeeds at the challenging task of identifying the points most vulnerable to membership inference attacks. The ability to achieve 60% precision at FPR=$10^{-3}$ is particularly notable given the difficulty of the task - we are identifying a small proportion of training points that can be confidently detected as members even under the most conservative attack settings that virtually eliminate false positives. The fact that we can match the performance of expensive SOTA shadow model-based approaches like Attack R, while using only training artifacts that are essentially free to collect, represents a significant advance in making privacy risk assessment more practical and accessible during model development. We believe that this high precision at strict FPR thresholds means model developers can reliably use LT-IQR to identify and address their models’ most severe privacy vulnerabilities during model development without the computational burden of training hundreds of shadow models.
5.4 Loss trace aggregation methods
We here evaluate different loss trace aggregation methods across multiple Precision@k and Recall@k thresholds (k = 1%, 3%, and 5% of the training data) to understand their effectiveness at identifying vulnerable points. Table 2 presents a comprehensive comparison of these methods.
The LT-IQR (75%-25%) method demonstrates consistently strong performance, achieving 61% Precision@1%. This suggests that the interquartile range of loss values robustly and effectively captures the pattern of vulnerability such as the ones we hypothesize in Section 4.1: points with high early-stage loss that eventually converge to low loss values are most likely to be vulnerable to membership inference attacks.
Mean loss emerges as the second-best performing aggregation method, achieving 36% Precision@1%. While this is significantly lower than LT-IQR, it still substantially outperforms simpler approaches like final loss (8%) and loss delta (9%). This suggests that considering the entire training trace provides more reliable signals of vulnerability than just examining the beginning and end states.
The relatively poor performance of loss delta (9% Precision@1%) and normalized loss delta (14% Precision@1%) is particularly interesting. These metrics attempt to directly capture our intuition about vulnerability being related to how much a point’s loss improves during training. However, they appear to be not robust enough to the noisy per-sample loss traces to effectively identify vulnerable points. This underscores the importance of considering the entire loss trace rather than just the total change in loss.
The final loss method performs poorly across all thresholds (8-12% precision), confirming our hypothesis that the final loss alone is insufficient for identifying vulnerable points. This poor performance can likely be attributed, as discussed in Section 4.1, to its inability to distinguish between naturally easy-to-classify points and those that required significant memorization to achieve low loss. We find it important to note the somewhat confusing difference between the final loss method and the LOSS attack. LOSS attack aims to distinguish members from non-members, and assigns higher probability of being a member to samples with low final loss. It stems from the intuition that memorized points would have low loss at the end of the training. Conversely, final loss makes the opposite hypothesis - among members, those with higher loss are more vulnerable. This highlights the difference in tasks those methods are designed for: the former is optimized for distinguishing members from non-members, while the latter identifies the most vulnerable points among members.
As we increase k from 1% to 5%, we observe a general trend of decreasing precision across all methods, which is expected as we include more borderline cases in our predictions. However, the relative ordering of method performance remains largely consistent, with LT-IQR maintaining its substantial lead over other aggregation approaches.
In terms of recall, we see similar patterns with LT-IQR achieving the highest Recall@1% (13%) among our proposed methods. This shows that our method is not only precise but also captures a significant portion of the truly vulnerable points, even when making a very selective prediction (top 1% of points).
This comparative analysis shows that while several aggregation methods can extract useful signals from loss traces, the LT-IQR approach provides the best balance of precision and recall, making it in our opinion a practical tool for identifying vulnerable points during model training.
5.5 Computational overhead
We empirically compute the computational overhead incurred by using our method and show that collecting losses during the training has a negligible impact on training time.
Table 3 presents the average time per epoch across 10 epochs when training a Wide ResNet28-2 [43] model on CIFAR-10 dataset on one A100 GPU. Our results show that collecting per-sample losses during training adds no statistically significant overhead. The baseline training time of $20.86\pm 1.02$ seconds per epoch and the time with loss collection ($21.08\pm 1.16$ seconds) have overlapping confidence intervals, confirming that tracking individual sample losses can be seamlessly integrated into the training pipeline with effectively zero cost. Even in cases where data augmentation requires an extra forward pass to compute losses on non-augmented samples, the overhead remains low. This worst-case scenario requires $24.84\pm 1.50$ seconds per epoch, representing roughly a 20% increase over the baseline training time.
These overheads are particularly negligible when compared to state-of-the-art membership inference attacks like LiRA and Attack R, which require training hundreds of shadow models. With LiRA requiring 256 shadow models and Attack R requiring 128 in our setup, their computational costs are respectively 256$\times$ and 128$\times$ that of the original training process. In contrast, our method’s maximum overhead of 1.2$\times$ makes it a practical solution for assessing privacy risks during model development.
Time per epoch (s)
Normal training $20.86\pm 1.02$
Normal training
+ collecting “free” losses $21.08\pm 1.16$
Normal training
+ extra forward pass $24.84\pm 1.50$
Table 3: Average time per epoch (s) across 10 epochs to train a Wide ResNet28-2 on CIFAR-10 on an A100 GPU.
5.6 Model-level vulnerability
Figure 6: Distribution of sample LT-IQR scores as computed on the partial loss traces at various epochs of training. Each line represents the LT-IQR distribution on the entire training dataset after a given particular epoch.
While our primary focus is on identifying individual vulnerable points, analyzing the distribution of vulnerability scores across training epochs provides insights into the overall model risk. Figure 6 visualizes the density distribution of LT-IQR scores across all training samples at different epochs, revealing several key patterns about model-level vulnerability to membership inference attacks.
Early in training (epochs 10-30), the LT-IQR distribution is relatively concentrated, with most samples showing similar levels of loss variation. This suggests that the model is still learning general patterns and has not yet begun significant memorization of individual samples. The tight distribution indicates that the model treats most samples similarly during this phase.
As training progresses (epochs 30-80), we observe that the distribution is shifting to the left, with more points having low LT-IQR. This makes remaining points with high LT-IQR scores more distinguishable and vulnerable. This transformation in the distribution shape corresponds to the model beginning to memorize individual samples, particularly outliers. Intuitively, the small proportion of points that are still experiencing higher loss volatility mid-training are the ones likely to become vulnerable to membership inference attacks.
By the final stages of training (epochs 80-100), the distribution takes on a power law-like shape, with a large concentration of samples having low LT-IQR values and a sharp but smooth decay toward higher values. This distribution shape aligns with previous findings about the non-uniform nature of privacy risk across training data [13, 5], where memorization tends to concentrate across a small subset of challenging examples.
The evolution of this distribution provides a useful signal about whether a model is vulnerable to membership inference attacks, and when it is most vulnerable. The transformation from a concentrated distribution to a power law-like shape suggests increased memorization and thus greater privacy risk. This observation could be valuable for model developers in deciding the risk posed by a model in practice but also when to halt training, apply privacy-preserving techniques like differential privacy, or selective unlearning of high-risk samples.
More specifically, the shape of the final LT-IQR distribution could serve as a quick diagnostic tool for assessing overall model privacy risk without requiring expensive shadow model training. The exponent of the power law could indicate how uniformly or non-uniformly privacy risk is distributed across the training samples.
5.7 Ablations
We conducted extensive ablation studies to understand how different parameter choices affect our method’s performance and to justify our design decisions. We focused on two key aspects:
  1. 1.
    The quantile parameters $q_{1}$ and $q_{2}$ used in the LT-IQR calculation
  2. 2.
    The value of k used in our evaluation metrics (Precision@k and Recall@k)
5.7.1 Impact of Quantile Parameters
The LT-IQR score relies on two quantile parameters $q_{1}$ and $q_{2}$ that determine which portions of the loss trace are used to compute the interquartile range. To thoroughly evaluate their impact, we conducted a grid search over:
  • $q_{1}$ values in [0.0, 0.50] at 0.05 intervals
  • $q_{2}$ values in [0.50, 1.00] at 0.05 intervals
We fixed FPR=$10^{-3}$ and k=1% of training data (250 points) for these experiments.
The results, visualized in Figure 7, reveal several important patterns.
Poor performance at extremes
Very low values of $q_{1}$ (near 0) and very high values of $q_{2}$ (near 1) lead to reduced performance. For the low values of $q_{1}$ this can be explained by early training epochs containing mostly initialization noise rather than meaningful learning patterns. For high $q_{2}$ values, we argue that by the final few epochs many points will have converged to near-zero losses. IQR would therefore fail to distinguish between traces with different convergence trajectory. As the final loss (a surrogate for ${q_{1}}=0.0$) can be influenced by both the example hardness, and memorization, two samples with different difficulty levels exhibiting different loss traces, can not be differentiated by the final loss alone.
Optimal middle range
The highest precision is achieved with $q_{1}$ around 0.3 and $q_{2}$ around 0.7. This suggests that the most informative part of the loss trace lies in the middle stages of training, where the model has moved past initial noise, but a large proportion of samples has not yet reached extremely low loss values.
Robustness
The performance is relatively stable within the middle ranges, showing that our method is not overly sensitive to the exact choice of quantiles. This is practically important as it means the method does not require careful tuning of these parameters.
Based on these findings, we selected $q_{1}=0.25$ and $q_{2}=0.75$ as our default parameters, as these values lie within the high-performance region and correspond to conventional quartile definitions.
5.7.2 Impact of k in Precision@k and Recall@k
We evaluate how the choice of k affects our method’s performance by testing $k$ values ranging from 0.1%-5% of the training data, corresponding to 25-1250 points respectively. Figure 8 presents Precision@k and Recall@k with varying values of $k$ for our best performing method LT-IQR.
At strict FPR thresholds ($\leq 10^{-3}$), where we focus on the most confidently believed to be identified vulnerable samples, we observe that smaller k values (0.1%-1%) achieve higher precision but at the cost of lower recall. This is true for both our method and other SOTA MIA methods and indicates that our highest-ranked samples are indeed likely to be vulnerable, but such selective prediction sets inevitably miss many samples believed to be vulnerable by another method. This limitation is particularly evident when using k=0.1%, which achieves the highest precision but can capture at most 2.5% of vulnerable samples, given that $|V_{10^{-3}}|=1,153$.
As we increase the FPR threshold beyond 0.01%, the performance hierarchy shifts significantly. Larger $k$ values (3%-5%) begin to outperform smaller ones in precision, as the set of vulnerable samples $V_{\alpha}$ grows with FPR. In this regime, the additional true positives identified with larger $k$ values outweigh the cost of including more false positives. However, we observe that recall decreases across all $k$ values as the vulnerable set expands, reflecting the increasing difficulty of capturing a larger set of vulnerable samples with a fixed-size prediction set.
These patterns reveal the inherent difficulty of the task and tension in selecting $k$. Small $k$ values excel at identifying the most obviously vulnerable samples with high confidence, while larger $k$ values provide better coverage but with lower confidence in individual predictions. Mid-range values around $k=1\%$ often provide a practical balance, identifying a meaningful number of vulnerable samples while maintaining reasonable precision.
The optimal choice of $k$ ultimately depends on the specific use case. For targeted privacy interventions where remediation is expensive, small $k$ values help focus resources on the most critical cases. Conversely, general privacy auditing might benefit from larger $k$ values to ensure comprehensive coverage. When specific requirements are not well-defined, $k=1\%$ serves as a reasonable default that balances these competing concerns.
Figure 7: Heatmap showing the Precision@1% for the LT-IQR method with varying quantile parameters of $q_{1}$ and $q_{2}$.
Figure 8: Precision@k (a) and Recall@k (b) of the LT-IQR method for varying values of k.
6 Conclusion
This work introduces a novel, computationally efficient approach to identify training samples most vulnerable to membership inference attacks which we call artifact-based.
Rather than attempting to classify all training samples as vulnerable or safe, we focus specifically on identifying the small subset of samples at highest privacy risk – a more challenging but crucial task for practical privacy protection. At a strict false positive rate of 0.001, these highly vulnerable samples make up less than 5% of the training data but represent the most serious privacy risks, as they have been shown to be confidently identifiable as training set members even under conservative attack settings.
Our method analyzes loss traces collected during model training to identify potentially vulnerable samples without requiring prohibitive shadow model training. Through extensive experiments on computer vision models and the CIFAR-10 dataset, we demonstrate that our Loss Trace IQR (LT-IQR) method achieves comparable performance to state-of-the-art shadow model-based approaches, while requiring only a fraction of the computational resources.
The key findings of our work include:
  • We analyze the multiplicity of at-risk samples, showing that state-of-the-art MIAs show only partial agreement on which samples are particularly vulnerable despite assuming the same attack and using similar methods.
  • Loss traces collected during training contain sufficient information to identify vulnerable samples, with our LT-IQR method achieving 61% precision at identifying the top 1% most vulnerable samples as determined by the LiRA attack, a performance in line with other shadow model-based approaches (Attack-R).
  • The computational overhead of collecting loss traces during training is negligible: near-zero in a baseline scenario, and less than $20\%$ in worst-case scenarios with data augmentation. This makes our method the most practical one for real-world deployment.
  • The evolution of loss trace distributions throughout training provides insights into how model memorization develops, potentially offering a new lens to evaluate model risks.
These results have important implications for privacy-aware machine learning development. The ability to identify vulnerable samples during training, without significant computational overhead and as part of standard toolboxes, enables model developers to assess privacy risks early and during model development and implement targeted mitigations such as selective unlearning or differential privacy. This is particularly valuable for large-scale models where training hundreds of shadow models would be prohibitively expensive.
7 Discussion
7.1 Limitations
While our method successfully identifies individual training samples vulnerable to membership inference attacks, some limitations and open questions remain.
Calibration Challenges
One notable limitation of our method is the difficulty in selecting appropriate threshold for vulnerability score in making a binary assessment. Threshold calibration is often side-stepped in the MIA literature [7, 33, 41], with most attacks evaluated with threshold-independent methods. This lack of a principled way to calibrate scores means that practitioners might have to rely on relative rankings rather than absolute measures of vulnerability.
Model-Level Vulnerability
Although our approach effectively identifies vulnerable individual samples, it does not readily provide a model-level vulnerability score. The distribution of LT-IQR scores provides insights into overall memorization patterns (as shown in Section 5.6) but more work is necessary to aggregate individual vulnerability scores into a model-level privacy risk metric. This is a limitation for practitioners who need to make binary decisions about whether a specific model is safe to release.
7.2 Future work
Test Set Analysis
So far in this paper we have only used the individual loss traces to predict vulnerability. The model trainer, however, additionally has access to dataset-level information, including loss traces of other training samples, as well as loss traces for the test samples, assuming evaluation is run regularly throughout training. Intuitively, loss traces for the test samples could act as an approximation for the denominator in the LiRA attack score. Where LiRA and Attack R rely on the loss distribution from the shadow models trained on datasets excluding the target sample to estimate example hardness, we expect that one could use the distribution of the target model losses on similar non-member samples.
Threshold Calibration Methods
Future research might want to explore ways to better calibrate vulnerability scores, potentially by creating hybrid approaches that combine lightweight artifact analysis with selective shadow model training, or exploring the relationship between loss traces and concrete privacy risks in real-world applications. Leveraging the test set can also be potentially used for threshold calibration.
7.3 Societal Impact Statement
The primary goal of our research is to enable model developers with low-cost privacy risk assessment. By identifying privacy leakage before deployment, developers can implement targeted mitigation strategies. While an adversary could theoretically attempt a similar approach, their effectiveness would be severely limited without access to the training artifacts that are available to model developers. Therefore, the defensive advantages of early privacy leakage detection and mitigation here strongly outweigh any potential risks from adversarial exploitation. Even in the unlikely scenario where a sophisticated adversary gains access to training artifacts, our approach ensures that privacy vulnerabilities can be identified and mitigated during development, substantially reducing the attack surface before models reach production.